Louis is 26 years old. Her daughter is 4 years old. In how many years will Louis be double her daughters age?
Answer by macston(5194)

.
L=Louis's age=26; D=Daughter's age=4
.
x=years until Louis's age is twice daughter's age
.
L+x=2(D+x)
26+x=2(4+x)
26+x=8+2x
18=x
ANSWER: Louis will be twice her daughter's age in 18 years.
.